OpenStudy (anonymous):

which or the binomials is a factor of this trinomial? x^2+3x-28

OpenStudy (anonymous):

\(x^2+3x-28=x^2-4x+7x-28=x(x-4)+7(x-4)=(x-4)(x+7)\)